OpenStudy (samigupta8):

The chord AC of parabola y^2=4x subtends 90° angle at points B and D on parabola.Of A,B,C, and D are represented by t1,t2,t3,t4 then

OpenStudy (samigupta8):

|t2t4-t1t3| =?

OpenStudy (samigupta8):

@ganeshie8

ganeshie8 (ganeshie8):

Could you draw the parabola and label the points/lengths ?

OpenStudy (samigupta8):

I just got two equations which are:- (t1+t2)(t2+t3)=-4 (t1+t4)(t3+t4)=-4

ganeshie8 (ganeshie8):

I don't understand the question; The way you have phrased it is very confusing

OpenStudy (samigupta8):

Instead of the word "Of" in the question it is "if".

ganeshie8 (ganeshie8):

Sorry I don't get why you want to label the points by two names. Something seems to be wrong

ganeshie8 (ganeshie8):

It would help if you take a screenshot of entire question and attach

OpenStudy (samigupta8):

t1,t2,t3,t4 are the parameters of A,B,C AND D.

ganeshie8 (ganeshie8):

Do you mean t1, t2, t3, t4 are the "position vectors" of the points A, B, C, D ?

OpenStudy (samigupta8):

Like the point A is (at1^2, 2at1) , we have assigned it a parametric term (t1) and likwise it is for the rest of them.

OpenStudy (samigupta8):

Did you get it?

ganeshie8 (ganeshie8):

y^2 = 4x parametric form would be (t^2, 2t)

ganeshie8 (ganeshie8):

Are you saying A = (t1^2, 2t1) B = (t2^2, 2t2) C = (t3^2, 2t3) D = (t4^2, 2t4) ?

OpenStudy (mww):

Yup that's how he assigned the paramaters.

OpenStudy (mww):

I'm guessing you would find the gradients of AB, AD, BC and CD

OpenStudy (samigupta8):

She* And yes ganeshie sir that's what i mean.

ganeshie8 (ganeshie8):

then you should end up with the equations (t1+t2)(t2+t3) = -1 (t1+t4)(t4+t3) = -1 right ?

OpenStudy (samigupta8):

Why so? We have a slope of any chord to be 2/t1+t2 for a parabola.

ganeshie8 (ganeshie8):

okay -4 on the right hand side lets see where to go from here

OpenStudy (samigupta8):

We didn't use the value of a of parabola in coming to this conclusion though.I think second equation must result from use of that.

OpenStudy (samigupta8):

@parthkohli

OpenStudy (samigupta8):

@thomas5267

OpenStudy (thomas5267):

I don't even understand the question. I am too tired now anyway.